OpenStudy (anonymous):

As a ball baring rolls along a straight track, it makes 11.0 revolutions while traveling a distance of 109mm. Find its radius.

OpenStudy (anonymous):

eleven times circumference = 109 mm ?

OpenStudy (anonymous):

11.0 times the idstance of 109mm

OpenStudy (anonymous):

11.0 revolutions, which means turns?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it is a straight path, so the ball is like a circle rolling

OpenStudy (anonymous):

11.0* pi*109?

OpenStudy (anonymous):

should be an equation 109 =

OpenStudy (anonymous):

11.0*pi

OpenStudy (anonymous):

34.6

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

whats the circumference of a circle

OpenStudy (anonymous):

3.14*diameter

OpenStudy (anonymous):

yes, and 2r = diameter. when using radius instead of diameter: 3.14 * 2 r

OpenStudy (anonymous):

eleven times the circumference of circle is 109 mm

OpenStudy (anonymous):

3.14*2*109=684.52

OpenStudy (anonymous):

109 mm = 11.0 *(3.14*2r)

OpenStudy (anonymous):

69.1

OpenStudy (anonymous):

yes: 109 mm = 69.1 r

OpenStudy (anonymous):

okay i have that equation

OpenStudy (anonymous):

and then divide 109 by 69.1 to get the radius

OpenStudy (anonymous):

in mm

OpenStudy (anonymous):

1.6 is the radius

OpenStudy (anonymous):

when rounded

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

good

OpenStudy (anonymous):

or could we do it like this.. 109/11.09=9.91=2*3.14*r

OpenStudy (anonymous):

r=9.91/2*3.14=15.56

OpenStudy (anonymous):

my computer is glitching to much 11.0

OpenStudy (anonymous):

looks good

OpenStudy (anonymous):

beware of decimal point i position

OpenStudy (anonymous):

so my final will be15.56 roundedto 15.6

OpenStudy (anonymous):

is this what you get?

OpenStudy (anonymous):

109/11.00 =9.91=2*3.14*r 4.955=3.14*r 1.58 = r

OpenStudy (anonymous):

don't worry about the rounding

OpenStudy (anonymous):

okay gotcha